OpenStudy (anonymous):

Joe has a collection of nickels and dimes that is worth $6.10. If the number of dimes was tripled and the number of nickels was increased by 49, the value of the coins would be $18.95. How many nickels and dimes does he have?

OpenStudy (anonymous):

What is 49 times the value of a nickel?

OpenStudy (anonymous):

^Yeah, I'm not too sure of the values either.

OpenStudy (anonymous):

i dont know im so confused

OpenStudy (anonymous):

If you figure out how much those 49 nickels are worth and add them to the original 6.10 then you can take the rest of the money and see how many dimes it is.

OpenStudy (anonymous):

It's 49 more nickels, not 49 times as many

OpenStudy (anonymous):

You have to multiply the 49 by .o5 because then you know how much the 49 nickels value @JakeV8

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

Confirmation on what @JakeV8 said. That 2nd equation should be (49 + x)(0.05) + 3y(0.10) = 18.95

OpenStudy (anonymous):

wait for @tcarroll010's answer... it will help :)

OpenStudy (anonymous):

you guys are confusing me haha

OpenStudy (anonymous):

After you get the value of the nickels, x, you add it to the original 6.10. What do you get? @jdorta1

OpenStudy (anonymous):

Just take Mathmuse's first equation and the second equation I gave, which JakeV8 actually gave.

OpenStudy (anonymous):

@tcarroll010 so its x(0.05)+y(0.10)=6.10 and (49+x)(0.05)+3y(0.10)=18.95?

OpenStudy (anonymous):

Yes, but credit goes to Mathmuse and JakeV8.

OpenStudy (anonymous):

ok so what do i do from there? i am seriously horrible at this stuff

OpenStudy (anonymous):

It might sound simplistic, but I'd get rid of those decimal points first by multiplying each side of both equations by 100.

OpenStudy (anonymous):

ok so it would be x(5)+y(10)=6.10?

OpenStudy (anonymous):

Perhaps you should listen to @Mathmuse or @JakeV8 at this point, because whether you listen to me or them, I don't deserve the credit here, but I'll stick around to help. I suggest you give one or the other a medal when we're done and then I'll medal the other guy. And that's 610 on the right side.

OpenStudy (anonymous):

OK

OpenStudy (anonymous):

I am having trouble keeping the connection going here... sorry... my page froze for the last 2 minutes or so. I don't mind helping either, and I'm not worried about the credit. I just hate making a comment and then having the page freeze, leaving everyone waiting :( I was going to medal tcarroll010 for helping :) Glad we're all friendly here :)

OpenStudy (anonymous):

@jdorta1, when you multiply that first equation through by 100, you need to be sure to multiply both sides... so the right side goes from 6.10 to 100*6.10 or 610.

OpenStudy (anonymous):

OK I GOT THAT

OpenStudy (anonymous):

so i now have x(5)+y(10)=610

OpenStudy (anonymous):

opps sorry for caps lol

OpenStudy (anonymous):

5(49 + x) + 30y = 1895 for the other equation. With this equation, expand 5(49 + x)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@jdorta1, What did you get for the 2nd equation after you expanded the stuff on the left like @tcarroll010 suggested?

OpenStudy (anonymous):

@JakeV8 oh my gosh i feel so dumb.. i have no idea what your talking about..

OpenStudy (anonymous):

Equation 1: x(5)+y(10)=610 Equation 2: 5(49 + x) + 30y = 1895 Now simplify equation 2 by multiplying the 5 across the parenthesis... expand by multiplying 5(49 + x). Then rewrite both Equation 1 and the new simplified version of Equation 2.

OpenStudy (anonymous):

@JakeV8 so it would be 245+5x+30y=1895?

OpenStudy (anonymous):

yes, good. Now, simplify more by moving the 245 from the left to the right of the equation... you have to subtract 245 from both sides. Then re-write the updated version of Equation 2.

OpenStudy (anonymous):

5x+30y=1651?

OpenStudy (anonymous):

check that subtraction again....

OpenStudy (anonymous):

oh shoot my bad haha its 1650

OpenStudy (anonymous):

ok, so you should have two equations with the x's and y's on the left, and the numbers on the right... that's the form you need to aim for when you're setting this up. Equation 1: 5x + 10y = 610 Equation 2: 5x + 30y = 1650

OpenStudy (anonymous):

Now you can finally get started on solving for x and y.

OpenStudy (anonymous):

@JakeV8 yes i have that

OpenStudy (anonymous):

@JakeV8 ok...how? lol

OpenStudy (anonymous):

@JakeV8 dont i subtract y from both sides or something

OpenStudy (anonymous):

Well, do you see how there is a 5x in both equations? The best technique here is called "elimination"....

OpenStudy (anonymous):

@JakeV8 ok i see that..

OpenStudy (anonymous):

Equation 1: 5x + 10y = 610 Equation 2: 5x + 30y = 1650 Let's subtract the first equation from the second. When you subtract, since "5x" appears in each, it will be eliminated. Eq 2: 5x + 30y = 1650 - Eq 1: 5x + 10y = 610 -------------------------

OpenStudy (anonymous):

@jakeV8 So its now 20y=1040

OpenStudy (anonymous):

@jakeV8 then i divide both sides by 20 correct?

OpenStudy (anonymous):

yay! Good thinking :)

OpenStudy (anonymous):

@JakeV8 y=52

OpenStudy (anonymous):

Correct! So, that means you've solved for "y". How will you solve for "x" now?

OpenStudy (anonymous):

@JakeV8 plug in y 5x+30(52)=1650

OpenStudy (anonymous):

You're doing well... continue :)

OpenStudy (anonymous):

5x+1560=1650 5x=90 x=18

OpenStudy (anonymous):

Correct again :) So, here's the place where you make sure you get this right... terrible to do all that work and then miss it... you need to double check your solutions in the original problem to make sure everything works. Joe has a collection of nickels and dimes that is worth $6.10. If the number of dimes was tripled and the number of nickels was increased by 49, the value of the coins would be $18.95. How many nickels and dimes does he have? You found x = 18 and y = 52. So you should use those numbers and see if they work in the word problem.

OpenStudy (anonymous):

@JakeV8 ok and how do you do that?

OpenStudy (anonymous):

18 nickels: 0.05 * 18 = $0.90 52 dimes: 0.10 * 52 = $5.20 ------------------------------ $6.10 <---- works so far...

OpenStudy (anonymous):

ooook I see where your going...

OpenStudy (anonymous):

The other statement said that if you tripled your dimes and had 49 nickels, the new value would be $18.95. 76 JakeV8 Medals 1 49 + 18 nickels = 67 nickels: 0.05 * 67 = $3.35 3 * 52 dimes = 156 dimes: 0.10 * 156 = $15.60 ------ -------------------------------------- $18.95 (this also works) So it's correct!! 18 nickels and 52 dimes.

OpenStudy (anonymous):

oops, some weird copy paste thing happened there with the medals stuff ??

OpenStudy (anonymous):

YAY!!! thank you so much you are the greatest!!! <3

OpenStudy (anonymous):

I think I must have copied my name and score or something... weird. oh well. Do you follow the "double-check" process? If it doesn't all work out, you know you made an error somewhere earlier. But if it does all work in both equations, then you are 100% guaranteed to get it correct :)

OpenStudy (anonymous):

you are awesome! lol

OpenStudy (anonymous):

Glad to help... :) once you are comfortable solving a system of 2 equations like this, they're all pretty similar. The other tricky thing is to learn to "translate" the word problem into math-language... you have to get those words into a form where it's 2 equations with 2 unknowns. The big clue is that they ask for "how many nickels and how many dimes?" so you know you have 2 unknowns... therefore you need 2 equations from somewhere... Good luck!

OpenStudy (anonymous):

I really do appreciate the help =)